Q3. On mountainous downhill roads, escape routes are sometimes placed on the side of the road for trucks whose brakes might fail. Assuming a constant upward slope of 26, calculate the horizontal and vertical components of the acceleration of a truck that slowed from 110 kms to rest in 7.0 seconds.
Answer #1

We know the equation of motion,

v = u + at ..................................(1)

Where, v is the final velocity of the object

u is the initial velocity of the object

a is acceleration of the object

t is time of motion

110 km/s 26

According to the question,

u = 110 km/s = 110 x 103 m/s

v = 0 m/s (as the truck finally stops)

t = 7.0 s

Putting these values in equation (1), we get

0 1 10 × 103 + a × 7.0

or,   7a-110 x 10

or,   a-一15714.3 m2/s2

or,   a15.7 km/s Cl

The negative sign is indicating that the acceleration is in opposite direction of velocity (i.e. in upward direction) .

a 15.7 km/s2 110 km/s 26

The components of the acceleration are

a 15.7 km/s* 0 26 a,

Where,

ar = a cos 26 = 15.7 × 0.8987 14.11 km/s

and a,, = a sin 26 = 15.7 × 0.4383 = 6.882 km/s2

Hence, the acceleration in component form can be written as

a = a_{x} , (- hat{i}) + a_{y} , hat{j} = -14.11 , , hat{i} + 6.882 , , hat{j}

Hence, the horizontal component of acceleration is 14.11 km/s2 in negative - x direction and the vertical component of acceleration is 6.882 km/s2 in y - direction (i.e. in upward direction).

For any doubt please comment and please give an up vote. Thank you.